The following table shows the production possibilities of two countries in making cars and airplanes:

 

Hours needed to make one:

Amount produced in 2400 hours:

 

Car

Airplanes

Cars

Airplanes

Japan

30

150

80

16

United States

50

200

48

12

a.     What is the opportunity cost of a car in Japan? What is the opportunity cost of an airplane in Japan? (2)

b.     What is the opportunity cost of a car in USA? What is the opportunity cost of an airplane in USA? (2)

c.     By looking at the table, say which city has an absolute advantage in both the goods? Why? (2)

Japan has a comparative advantage in which good and USA has a comparative advantage in which of the two goods? Why? 


1
Expert's answer
2022-01-23T15:40:07-0500

Solution:

a.). Opportunity cost refers to the next best-foregone alternative.

The opportunity cost of a car in Japan is = "\frac{16}{48} = \frac{1}{3}" airplane.


The opportunity cost of an airplane in Japan is = "\frac{48}{16}" = 3 cars

 

b.). The opportunity cost of a car in USA is = "\frac{15}{60} =\frac{1}{4}" airplane.


The opportunity cost of an airplane in the USA is = "\frac{60}{15}" 60/15 = 4 cars

 

c.). The USA has an absolute advantage in both of the goods. This is because it has the ability to produce more of the two products than Japan.

 

Japan has a comparative advantage in producing airplanes. This is because it has the ability to produce airplanes at a lower opportunity cost than the USA. It produces more airplanes compared to the USA using fewer hours.

The USA has a comparative advantage in producing cars. This is because it has the ability to produce cars at a lower opportunity cost than Japan. It produces more cars using fewer hours.